OpenStudy (anonymous):

A juggler throws balls into the air. He throws one when the previous one is at its highest point. How high the balls rise if he throws m balls per second?

OpenStudy (anonymous):

Use s = ut + 1/2 at^2, t = 1/m, u = 0, a = 9.81